OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -2 - 6t. Find the instantaneous velocity at t = 2 by finding the derivative.

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

as I remember it, the velocity is s'(t) what is the first derivative of -2-6t

OpenStudy (anonymous):

(d/dx)(-2-6t)?

OpenStudy (bibby):

well you just rephrased my question

OpenStudy (phi):

it would be d/dt (-2 - 6t) which is (term by term \[ \frac{d}{dt} (-2) -6 \frac{d}{dt}t \]

OpenStudy (phi):

constants don't change with time. in other words d/dt (-2) is 0 that leaves \[ -6 \frac{d}{dt} t \]

OpenStudy (anonymous):

okay, so what do I do with the derivative?

OpenStudy (anonymous):

@bibby @phi

OpenStudy (bibby):

take the derivative of t

OpenStudy (anonymous):

-12t?

OpenStudy (phi):

after you get s'(t) , which in this case is s'(t)= -6 (a constant) you evaluate it at t=2 you (of course) get -6

OpenStudy (anonymous):

or just -12?

OpenStudy (phi):

if you don't know how to find the derivative of d/dt t then you need to review the intro...

OpenStudy (bibby):

\(\frac{d}{dt}(-2-6t)=\frac{d}{dt})-2_-\frac{d}{dt}6t\)

OpenStudy (anonymous):

I thought I had to plug in 2 for t though

OpenStudy (phi):

you do, after you find s'(t) the steps are: s(t) = -2 - 6t you take the derivative with respect to "t" d/dt (-2) is 0 d/dt (-6t ) is -6 you get s'(t)= -6 now you find s'(2) but s'(2) = -6 (in other words, the "function" s'(t) is a constant value, always equal to -6)

OpenStudy (phi):

because s'(t) does not have a "t" in it... you just get -6 no matter what t is

OpenStudy (anonymous):

so s'(t)=-6 with or without the t?

OpenStudy (phi):

I would say: so s'(t)=-6 there is no "t" it sometimes happens. But it's ok, (it's like y= -6, a straight line, with y always = -6 no matter what x is... same idea here)

OpenStudy (anonymous):

so what's the instantaneous velocity then?

OpenStudy (phi):

s'(t) evaluated at t=2

OpenStudy (phi):

and because s'(t) is constant , always equal to -6 s'(2) is -6 i.e. instantaneous velocity is -6

OpenStudy (anonymous):

okay, and just for my own record because this wasn't in my lesson at all, what exactly is an instantaneous velocity? I know velocity is speed with direction but idk about the instantaneous velocity...

OpenStudy (phi):

velocity is change in distance divided by change in time (sounds vaguely like the definition of slope, if distance is y and time is x) if the velocity is always changing (speeding up or slowing down) then to find velocity we have to look at the change in distance over a *very small change in time* if we plot distance over time, and velocity is changing, the curve would not be a straight line. The slope (velocity) can only be approximated as the slope of a tangent line to the curve.

OpenStudy (anonymous):

Okay, thanks :D

OpenStudy (phi):

If you have time, watch Khan's videos.